OpenStudy (anonymous):

Solve x2 - 8x - 20 = 0. x = 4, x = -5 x = -4, x = 5 x = 2, x = -10 x = -2, x = 10

OpenStudy (lucaz):

look for two numbers the product is -20 and the sum is -8 try -10 and 2 (-10)(2)= -20 and -10+2=-8 x^2-8x-20=(x-10)(x+2), the roots are x=10 and x=-2

OpenStudy (anonymous):

Thank you very much @lucaz !

OpenStudy (lucaz):

welcome